# Survey on Factuality in Large Language Models: Knowledge, Retrieval and Domain-Specificity

### v1

@@ 5.1.5 Decoding
5.2 On Retrieval-Augmented Generation
@@
(2) Interactive Retrieval and Generation (Sec 5.2.2): Examples here include the integration of Chain-of-Thoughts steps into query retrieval (He et al., 2022) and the use of an LLM-based agent framework that taps into external knowledge APIs (Yao et al., 2023a).
(3) Adapting LLMs to the RAG Setting (Sec 5.2.3): This involves methods like the one proposed by Peng et al. (2023), which combines a fixed LLM with a plug-and-play retrieval module. Another notable approach is REPLUG (Shi et al., 2023), a retrieval-augmented framework that treats the LLM as a black box and fine-tunes retrieval models using language modeling scores.
(4) Retrieving from Additional Knowledge Bases (Sec 5.2.5 and Sec 5.2.4): This category includes methods that retrieve from external parametric memories (Chen et al., 2023a) or knowledge graphs (Zhang et al., 2023d) to enhance the model’s knowledge base.
5.2.1 Normal RAG Setting
@@
Lazaridou et al. (2022) present a method that capitalizes on the few-shot capabilities of large-scale language models to enhance their grounding in factual and current information. Drawing from semi-parametric language models, the approach conditions LMs using few-shot prompts based on data sourced from Google Search. For any given query, the method retrieves pertinent documents from the web, extracts the top 20 URLs, and processes them to obtain clear text. These documents are segmented into paragraphs, and the most relevant ones are chosen using TF-IDF based on their similarity to the query. The LMs are then conditioned using few-shot prompts that incorporate the retrieved paragraphs. This k-shot prompting technique is augmented with an evidence paragraph, creating a prompt structure that encompasses evidence, query, and response. The method also involves generating multiple answers from the model and reranking them using different probabilistic factorizations. The experimental results indicate that by conditioning on retrieved evidence, the 7B Gopher LM (Rae et al., 2022a) surpassed the performance of the 280B Gopher LM, with relative improvements reaching up to 30% on the NQ (Kwiatkowski et al., 2019) dataset.
5.2.2 Interactive Retrieval
@@
In recent studies, there is a growing interest in integrating Chain-of-Thoughts (Wei et al., 2022) steps into query retrieval. He et al. (2022) introduce a method that generates multiple reasoning paths and their corresponding predictions for each query. This process involves retrieving relevant knowledge from external sources like Wikidata (Vrandečić and Krötzsch, 2014), WordNet (Miller, 1992), and ConceptNet (Speer et al., 2017). The faithfulness of each reasoning path determines based on entailment scores, contradiction scores, and MPNet similarities (Song et al., 2020) with the retrieved knowledge. The prediction with the highest faithfulness score is chosen as the final result. This method demonstrates superior performance in tasks such as commonsense reasoning (Geva et al., 2021b), temporal reasoning (Jia et al., 2018), and tabular reasoning (Gupta et al., 2020b), outperforming the baseline CoT reasoning and self-consistency methods (Wang et al., 2023e). On a related note, Trivedi et al. (2023) propose IRCoT, an innovative retrieval technique that interweaves the CoT process. In this approach, each generated sentence during the CoT combines with the question to form a retrieval query. The subsequent reasoning step then produces by the Language Model using both the retrieval results and the prior reasoning. This interleaving method finds to enhance the performance of both retrieval and CoT in Open-domain QA. Experiments proves that this is beneficial for models of different sizes, including GPT-3(175B) (Brown et al., 2020) and Flan-T5 (Chung et al., 2022) families.
FLARE (Jiang et al., 2023b) is a dynamic solution to address the limitations of previous RAG works which either retrieve only once at the onset of generation or do so based on fixed intervals. Single retrievals are insufficient for long-form generation due to the evolving information needs during the process, and fixed intervals for previous token queries can be inappropriate, FLARE determines ”when and what to retrieve” dynamically. The decision of ”when” is based on whether the current sentence contains a token with a generation probability below a set threshold. If it doesn’t, the sentence is accepted and moves to the next generation step; otherwise, retrieval augmented generation occurs. For the ”what”, the current sentence is used as a query. To address the challenge of low-probability tokens affecting retrieval accuracy, two solutions are proposed: masking low-probability tokens, and using LLM for generation of these tokens as queries. Testing on tasks like Multihop QA (Ho et al., 2020), Commonsense Reasoning (Geva et al., 2021a), Long-form QA (Stelmakh et al., 2022), and Open-domain Summarization (Hayashi et al., 2021) using GPT3.5, results show FLARE outperforms baselines, with both query generation methods showing comparable performance.
@@
Yao et al. (2023a) present a new framework named ReAct that integrates Chain-of-Thoughts reasoning with action. Through in-context learning, the LLM’s CoT output is transformed into descriptions of reasoning processes and action behaviors. Subsequently, these action descriptions are standardized and executed, with the results being incorporated into the next prompt. In terms of results, for tasks like Fact checking and QA, while CoT has 14% of its correct answers containing incorrect reasoning steps or facts, ReAct only has 6%. In the incorrect answers, 56% of CoT’s errors were due to errors in reasoning steps or facts, whereas ReAct has no factual errors. However, ReAct have 23% of its errors resulting from search errors.
Shinn et al. (2023) propose a prompt engineering framework named Reflextion to enable LLMs to reflect on and correct previous errors. They use linguistic feedback to strengthen an agent’s actions instead of adjusting model weights. Specifically, the Reflextion Agent, an LLM, first interacts with its environment by generating an ”Action” via ReAct (Yao et al., 2023a) or Chain-og-Thoughts (Wei et al., 2022) in few-shot scenarios, which results in an ”Observation”. This ”Observation”, whether a reward, error message, or natural language feedback, provides insights on the Agent’s current ”Action”. When the Agent receives a failure signal, it triggers a self-reflextion mechanism, utilizing the LLM, to summarize the reasons for the failure into its Memory module, creating a ”long-term memory”. On subsequent generations, the Agent can review all past reflextion memories to prevent mistakes. Experimental findings indicate that Reflextion achieves a 10-20% performance increase over the baseline methods ReACT and CoT on datasets like AlfWorld (Shridhar et al., 2021), HotPotQA (Yang et al., 2018), and HumanEval (Chen et al., 2021a).
@@
5.2.3 Retrieval Adaptation
Recent research (Wang et al., 2023a; Ren et al., 2023) has highlighted that merely using the retrieved information in LLMs doesn’t always enhance their ability to answer factual questions. This underscores the importance of enabling LLMs to better adapt to the retrieved data to produce more accurate content. In this section, we delve into various strategies that facilitate this adaptation. Specifically, we explore three methodological approaches: prompt-based methods, SFT-based methods, and RLHF-based methods.
@@ 5.2.3 Retrieval Adaptation
Peng et al. (2023) introduce the LLM-Augmenter system, a system using a fixed LLM combined with a plug-and-play retrieval module to help the LLM perform better in tasks that are particularly sensitive to factual errors. This system enhances its performance by enabling the LLM to use a series of modules (e.g. allowing the LLM to interact with external knowledge) to assist the LLM in generating results grounded in evidence. And they use automated feedback generated by utility functions (e.g., the factuality score of a LLM-generated response) to modify LLM’s candidate response options. The author evaluates the system’s performance on information-seeking dialog (Galley et al., 2019b) and Wiki QA (Chen et al., 2021b) and experiments show that the system can significantly reduce ChatGPT’s errors without sacrificing the fluency and informativeness of the generated content.
@@
Shi et al. (2023) introduce REPLUG, a retrieval-augmented framework that considers the LLM as a black box, freezes its parameters and tunes retrieval models with supervision signals using language modeling scores. In this framework, the input context and the document are encoded through the dual encoder architecture, cosine similarity is then calculated to retrieve related documents. The likelihood for each retrieved document and the language model scores are computed. Then they can update the retrieval model parameters by minimizing the KL divergence between retrieved document likelihood and the language model”s score distribution. Ablation experiments demonstrate that this method significantly improves the performance of the original language models and the improvements are not coming from ensembling random documents.
@@
5.2.4 Retrieval on External Memory
@@
Li et al. (2022c) store knowledge in the form of key-value pairs in memory. The key is obtained by encoding knowledge using a Doc Retrieval Embedder, while the value is encoded using a Transformer encoder. Similar to traditional retrieval-based LLMs, the model encodes the input using a Query retrieval embedder and retrieves knowledge from memory. The retrieved value is then integrated into the model’s multi-head attention layer through cross-attention to enhance factuality.
@@
Similar to the three methods mentioned above, knowledge plugin (Zhang et al., 2023g) also introduce entity-related knowledge. However, instead of integrating the knowledge directly into the model layers, they utilize a pre-trained mapping network. This network maps the entity embeddings to the token embedding space of the Pre-trained Language Model (PLM). Ultimately, the mapped entity embeddings are injected at the input embedding level, facilitating the knowledge insertion process.
@@
Chen et al. (2023a) is a follow-up work to (Gao et al., 2023a) and (Thorne and Vlachos, 2021). Similar to EFEC, this paper fine-tunes a T5 model to serve as an editor, but it introduces negative samples during fine-tuning. The model PURR is trained to take user questions, perform Google searches to retrieve the top 5 web page summaries (used as positive samples), and generate noise by replacing some content in these positive samples using the language model. A sequence-to-sequence model is then trained to correct the noisy sentences back to their correct versions. This approach differs from EFEC, which used a mask-and-fill approach. PURR represents an improvement over EFEC, focusing on directly training a language model to edit incorrect sentences into correct ones using Google search for generating positive samples, ultimately leading to increased F1 scores.
5.2.5 Retrieval on Structured Knowledge Source
@@
Zhang et al. (2023d) utilize knowledge graphs (KG) for retrieval to tackle factual errors. They observe that there can be inconsistencies between a user’s request and the content in the KG. For instance, when a user mentions a full name, the KG might only have its abbreviation, leading to imperfect retrieval results. To rectify this, they propose a method to rephrase the user’s request. Their approach involves generating an SQL query based on the user’s input and the database metadata using a LLM. And then they query the database and ask the LLM to identify which entity in the sentence corresponds to an entity in the database, thereby creating a mapping. Using the entity names from the database, the LLM is prompted to reformulate the question. If a database query results in multiple rows for a selected column and item, a new question is generated using a greedy approach, prompting the user for more specific details until a conclusive answer is reached. Experiments show that this method exhibits notable enhancements in comparison to contemporary state-of-the-art techniques in mitigating language model inaccuracies.
StructGPT (Jiang et al., 2023a) is a general prompt framework to support LLMs reasoning on structured data (e.g., KG, Table, and Database). In general, the core of this framework is that they construct the specialized interfaces to collect relevant evidence from structured data (i.e., reading), and let LLMs concentrate on the reasoning task based on the collected information (i.e., reasoning). Specially, they propose an invoking-linearization-generation procedure to support LLMs in reasoning on the structured data with the help of the interfaces. By iterating this procedure with provided interfaces, our approach can gradually approach the target answers to a given query. Experiments conducted on three types of structured data, including KGQA, TableQA, and Text-to-SQL, show that StructGPT greatly improves the performance of LLMs, under the few-shot or zero-shot settings.
5.3 Domain Factuality Enhanced LLMs
@@
TABLE VIII: LLMs Enhanced for Domain-Specific Factuality. In the ‘domain’ column, we utilize the following abbreviations: healthcare/medicine (H), finance (F), law/legal (L), geoscience/environment (G), education (E), food testing (FT), and home renovation (HR).
Reference
Do-
main
Model
Eval
Task
Eval
Dataset
Continual
Pretrained?
Continual
SFT?
Train From
Scratch?
External
Knowledge
Zhang et al. (2023a) H
Baichuan-7B,
Ziya-LLaMA-13B
QA
cMedQA2,
WebMedQA,
Huatuo-26M
Yang et al. (2023c) H Ziya-LLaMA-13B QA
CMtMedQA,
huatuo-26M
Wang et al. (2023f) H
GPT-3.5-Turbo,
LLaMA-2-13B
QA
MedQAUSMLE,
MedQAMCMLE,
MedMCQA
Ross et al. (2022) H MOLFORMER
Molecule
properties
prediction
Bao et al. (2023) H Baichuan-13B QA
CMB-Clin,
CMD,
CMID
Guan et al. (2023) H ChatGPT
IU-RR,
MIMIC-CXR
Liu et al. (2023e) H GPT-4
Medical
Text
De-Identification
Li et al. (2023f) H LLaMA QA
Venigalla et al. (2022) H GPT (2.7b) QA
Xiong et al. (2023b) H ChatGLM-6B QA
Tan et al. (2023a) H Baichuan-7B QA C-Eval, MMLU
Luo et al. (2022) H GPT-2 QA, DC, RE
Jin et al. (2023) H Codex QA GeneTuring
Zakka et al. (2023) H text-davinci-003 QA ClinicalQA
Liu et al. (2023d) H GPT-2medium
Molecular
Property
Prediction,
Molecule-text
translation
Nguyen (2023) L GPT3
Savelka et al. (2023) L GPT-4
Huang et al. (2023a) L LLaMA CN Legal Tasks
Cui et al. (2023a) L Ziya-LLaMA-13B QA
national judicial
examination
question
Li et al. (2023e) F BLOOMZ
4 major tasks
12 subtasks
EcomInstruct
Wu et al. (2023) F BLOOM
Financial NLP
(SA, BC,
NER,
NER+NED, QA)
Financial
Datasets
Deng et al. (2023) G LLaMA-7B GeoBench
Bai et al. (2023) G ChatGLM-6B
Fan et al. (2023) E phoenix-inst-chat-7b
Chinese
Grammatical Error
Correction
ChatGPT-
generated,
Human-
annotated
Qi et al. (2023) FT Chinese-LLaMA2-13B QA
Wen et al. (2023) HR Baichuan-13B
C-Eval,
CMMLU,
EvalHome
Table VIII lists the domain-factuality enhanced LLMs. Here, we include several domains, including healthcare/medicine (H), finance (F), law/legal (L), geoscience/environment (G), education (E), food testing (FT), and home renovation (HR).
@@
(2) Continual SFT: Another strategy for enhancing the factuality of AI models. In this approach, the model is fine-tuned using labeled or annotated data specific to the domain of interest. This fine-tuning process allows the model to learn and adapt to the nuances and specifics of the domain, improving its ability to provide accurate and contextually relevant information. It can be particularly useful in applications where access to domain-specific labeled data is available over time, such as in the case of legal databases, medical records, or financial reports (Bao et al., 2023; Li et al., 2023f).
@@
For each Domain-specific LLM, we list its respective enhancement methods, which are presented in Table VIII.
5.3.1 Healthcare domain-enhanced LLMs
These LLMs have emerged as powerful tools in the medical field, offering a diverse range of capabilities. These models, such as CohortGPT(Guan et al., 2023), ChatDoctor(Li et al., 2023f), DeID-GPTLiu et al., 2023e, BioMedLMVenigalla et al., 2022, DoctorGLMXiong et al., 2023a, MedChatZHTan et al., 2023a, BioGPT(Luo et al., 2022), GeneGPT(Jin et al., 2023), Almanac(Zakka et al., 2023), and MolXPT(Liu et al., 2023d), harness the potential of LLMs to revolutionize healthcare. They are equipped with features like classifying unstructured medical text into disease labels, improving performance with knowledge graphs and sample selection strategies, fine-tuning on large datasets of patient-doctor dialogues, enabling automatic medical text de-identification, excelling in medical question-answering tasks, handling traditional Chinese medical question-answering, and outperforming in various biomedical NLP tasks. Some models interact with web APIs for genomics questions, while others specialize in clinical guidelines and treatment recommendations. These LLMs not only demonstrate state-of-the-art performance on healthcare domain but also emphasize the importance of domain-specific training and evaluation, showcasing their potential in transforming healthcare and clinical decision-making.

### v3

@@ 5.1.5. Decoding
5.2. On Retrieval-Augmented Generation
@@
(2) Interactive Retrieval and Generation (Sec 5.2.2): Examples here include the integration of Chain-of-Thoughts steps into query retrieval (He et al., 2022) and the use of an LLM-based agent framework that taps into external knowledge APIs (Yao et al., 2023b).
(3) Adapting LLMs to the RAG Setting (Sec 5.2.3): This involves methods like the one proposed by Peng et al. (2023), which combines a fixed LLM with a plug-and-play retrieval module. Another notable approach is REPLUG (Shi et al., 2023b), a retrieval-augmented framework that treats the LLM as a black box and fine-tunes retrieval models using language modeling scores.
(4) Retrieving from Additional Knowledge Bases (Sec 5.2.5 and Sec 5.2.4): This category includes methods that retrieve from external parametric memories (Chen et al., 2023e) or knowledge graphs (Zhang et al., 2023f) to enhance the model’s knowledge base.
5.2.1. Normal RAG Setting
@@
Lazaridou et al. (2022) present a method that capitalizes on the few-shot capabilities of large-scale language models to enhance their grounding in factual and current information. Drawing from semi-parametric language models, the approach conditions LMs using few-shot prompts based on data sourced from Google Search. For any given query, the method retrieves pertinent documents from the web, extracts the top 20 URLs, and processes them to obtain clear text. These documents are segmented into paragraphs, and the most relevant ones are chosen using TF-IDF based on their similarity to the query. The LMs are then conditioned using few-shot prompts that incorporate the retrieved paragraphs. This k-shot prompting technique is augmented with an evidence paragraph, creating a prompt structure that encompasses evidence, query, and response. The method also involves generating multiple answers from the model and reranking them using different probabilistic factorizations. The experimental results indicate that by conditioning on retrieved evidence, the 7B Gopher LM (et al., [n. d.]) surpassed the performance of the 280B Gopher LM, with relative improvements reaching up to 30% on the NQ (Kwiatkowski et al., 2019) dataset.
5.2.2. Interactive Retrieval
@@
In recent studies, there is a growing interest in integrating Chain-of-Thoughts (Wei et al., 2022) steps into query retrieval. He et al. (2022) introduce a method that generates multiple reasoning paths and their corresponding predictions for each query. This process involves retrieving relevant knowledge from external sources like Wikidata (Vrandečić and Krötzsch, 2014), WordNet (Miller, 1992), and ConceptNet (Speer et al., 2017). The faithfulness of each reasoning path determines based on entailment scores, contradiction scores, and MPNet similarities (Song et al., 2020) with the retrieved knowledge. The prediction with the highest faithfulness score is chosen as the final result. This method demonstrates superior performance in tasks such as commonsense reasoning (Geva et al., 2021b), temporal reasoning (Jia et al., 2018), and tabular reasoning (Gupta et al., 2020b), outperforming the baseline CoT reasoning and self-consistency methods (Wang et al., 2023g). On a related note, Trivedi et al. (2023) propose IRCoT, an innovative retrieval technique that interweaves the CoT process. In this approach, each generated sentence during the CoT combines with the question to form a retrieval query. The subsequent reasoning step then produces by the Language Model using both the retrieval results and the prior reasoning. This interleaving method finds to enhance the performance of both retrieval and CoT in Open-domain QA. Experiments proves that this is beneficial for models of different sizes, including GPT-3(175B) (Brown et al., 2020) and Flan-T5 (Chung et al., 2022) families.
FLARE (Jiang et al., 2023a) is a dynamic solution to address the limitations of previous RAG works which either retrieve only once at the onset of generation or do so based on fixed intervals. Single retrievals are insufficient for long-form generation due to the evolving information needs during the process, and fixed intervals for previous token queries can be inappropriate, FLARE determines ”when and what to retrieve” dynamically. The decision of ”when” is based on whether the current sentence contains a token with a generation probability below a set threshold. If it doesn’t, the sentence is accepted and moves to the next generation step; otherwise, retrieval augmented generation occurs. For the ”what”, the current sentence is used as a query. To address the challenge of low-probability tokens affecting retrieval accuracy, two solutions are proposed: masking low-probability tokens, and using LLM for generation of these tokens as queries. Testing on tasks like Multihop QA (Ho et al., 2020), Commonsense Reasoning (Geva et al., 2021a), Long-form QA (Stelmakh et al., 2022), and Open-domain Summarization (Hayashi et al., 2021) using GPT3.5, results show FLARE outperforms baselines, with both query generation methods showing comparable performance.
@@
Yao et al. (2023b) present a new framework named ReAct that integrates Chain-of-Thoughts reasoning with action. Through in-context learning, the LLM’s CoT output is transformed into descriptions of reasoning processes and action behaviors. Subsequently, these action descriptions are standardized and executed, with the results being incorporated into the next prompt. In terms of results, for tasks like Fact checking and QA, while CoT has 14% of its correct answers containing incorrect reasoning steps or facts, ReAct only has 6%. In the incorrect answers, 56% of CoT’s errors were due to errors in reasoning steps or facts, whereas ReAct has no factual errors. However, ReAct have 23% of its errors resulting from search errors.
Shinn et al. (2023) propose a prompt engineering framework named Reflextion to enable LLMs to reflect on and correct previous errors. They use linguistic feedback to strengthen an agent’s actions instead of adjusting model weights. Specifically, the Reflextion Agent, an LLM, first interacts with its environment by generating an ”Action” via ReAct (Yao et al., 2023b) or Chain-og-Thoughts (Wei et al., 2022) in few-shot scenarios, which results in an ”Observation”. This ”Observation”, whether a reward, error message, or natural language feedback, provides insights on the Agent’s current ”Action”. When the Agent receives a failure signal, it triggers a self-reflextion mechanism, utilizing the LLM, to summarize the reasons for the failure into its Memory module, creating a ”long-term memory”. On subsequent generations, the Agent can review all past reflextion memories to prevent mistakes. Experimental findings indicate that Reflextion achieves a 10-20% performance increase over the baseline methods ReACT and CoT on datasets like AlfWorld (Shridhar et al., 2021), HotPotQA (Yang et al., 2018), and HumanEval (Chen et al., 2021b).
@@
Self-RAG (Asai et al., 2023) builds upon the Retrieval-Augmented Generation (RAG) framework by incorporating a self-reflective approach. This includes on-demand retrieval, prompting the LLM to consider whether the retrieved documents are relevant and supportive of the argument, thereby enhancing the factuality of the LLM’s output. Self-RAG achieves this by having the model produce special tokens (i.e., reflection tokens) during the output process. The authors employ an end-to-end training method that enables the model to generate these special tokens. At the inference level, Self-RAG decodes a retrieval token for each input and each segment of the previously generated content. If the token is ’no,’ the model proceeds with the normal output; if ’yes,’ it performs a retrieval. The input, previous output, and each retrieved document are then fed into another session of the model, which assesses the relevance of each document. If a document is relevant, the model further evaluates whether it is supportive and should be used in generation. Based on this document, the model produces a segment and applies soft and hard constraints using the generated reflection token to select the most appropriate document. Self-RAG then incorporates the most suitable document into the continued generation of content, providing citations as needed. This process repeats until the entire content is generated. Self-RAG requires training data, which is annotated by GPT-4. Experiments based on the Llama-2 model on Short-form QA (PopQA (Mallen et al., 2023), TQA (Joshi et al., 2017)), Closed-set QA (Pub (Zhang et al., 2023e), ARC (Clark et al., 2018)), and Long-form NLG with Citations (Bio Generation (Min et al., 2023), ALCE-ASQA (Gao et al., 2023b)) have demonstrated its effectiveness.
5.2.3. Retrieval Adaptation
Recent research (Wang et al., 2023b; Ren et al., 2023) has highlighted that merely using the retrieved information in LLMs doesn’t always enhance their ability to answer factual questions. This underscores the importance of enabling LLMs to better adapt to the retrieved data to produce more accurate content. In this section, we delve into various strategies that facilitate this adaptation. Specifically, we explore three methodological approaches: prompt-based methods, SFT-based methods, and RLHF-based methods.
@@ 5.2.3. Retrieval Adaptation
Peng et al. (2023) introduce the LLM-Augmenter system, a system using a fixed LLM combined with a plug-and-play retrieval module to help the LLM perform better in tasks that are particularly sensitive to factual errors. This system enhances its performance by enabling the LLM to use a series of modules (e.g. allowing the LLM to interact with external knowledge) to assist the LLM in generating results grounded in evidence. And they use automated feedback generated by utility functions (e.g., the factuality score of a LLM-generated response) to modify LLM’s candidate response options. The author evaluates the system’s performance on information-seeking dialog (Galley et al., 2019b) and Wiki QA (Chen et al., 2021a) and experiments show that the system can significantly reduce ChatGPT’s errors without sacrificing the fluency and informativeness of the generated content.
@@
Shi et al. (2023b) introduce REPLUG, a retrieval-augmented framework that considers the LLM as a black box, freezes its parameters and tunes retrieval models with supervision signals using language modeling scores. In this framework, the input context and the document are encoded through the dual encoder architecture, cosine similarity is then calculated to retrieve related documents. The likelihood for each retrieved document and the language model scores are computed. Then they can update the retrieval model parameters by minimizing the KL divergence between retrieved document likelihood and the language model”s score distribution. Ablation experiments demonstrate that this method significantly improves the performance of the original language models and the improvements are not coming from ensembling random documents.
@@
5.2.4. Retrieval on External Memory
@@
Li et al. (2022a) store knowledge in the form of key-value pairs in memory. The key is obtained by encoding knowledge using a Doc Retrieval Embedder, while the value is encoded using a Transformer encoder. Similar to traditional retrieval-based LLMs, the model encodes the input using a Query retrieval embedder and retrieves knowledge from memory. The retrieved value is then integrated into the model’s multi-head attention layer through cross-attention to enhance factuality.
@@
Similar to the three methods mentioned above, knowledge plugin (Zhang et al., 2023h) also introduce entity-related knowledge. However, instead of integrating the knowledge directly into the model layers, they utilize a pre-trained mapping network. This network maps the entity embeddings to the token embedding space of the Pre-trained Language Model (PLM). Ultimately, the mapped entity embeddings are injected at the input embedding level, facilitating the knowledge insertion process.
Xue et al. (2023) address the challenge of improving factual consistency in knowledge-grounded dialogue systems by introducing extended feed-forward networks (FFNs) and leveraging reinforcement learning, resulting in more accurate and reliable responses, as demonstrated on the WoW (Dinan et al., 2019) and CMU_DoG (Zhou et al., 2018) datasets.
@@
Chen et al. (2023e) is a follow-up work to (Gao et al., 2023a) and (Thorne and Vlachos, 2021). Similar to EFEC, this paper fine-tunes a T5 model to serve as an editor, but it introduces negative samples during fine-tuning. The model PURR is trained to take user questions, perform Google searches to retrieve the top 5 web page summaries (used as positive samples), and generate noise by replacing some content in these positive samples using the language model. A sequence-to-sequence model is then trained to correct the noisy sentences back to their correct versions. This approach differs from EFEC, which used a mask-and-fill approach. PURR represents an improvement over EFEC, focusing on directly training a language model to edit incorrect sentences into correct ones using Google search for generating positive samples, ultimately leading to increased F1 scores.
5.2.5. Retrieval on Structured Knowledge Source
@@
Zhang et al. (2023f) utilize knowledge graphs (KG) for retrieval to tackle factual errors. They observe that there can be inconsistencies between a user’s request and the content in the KG. For instance, when a user mentions a full name, the KG might only have its abbreviation, leading to imperfect retrieval results. To rectify this, they propose a method to rephrase the user’s request. Their approach involves generating an SQL query based on the user’s input and the database metadata using a LLM. And then they query the database and ask the LLM to identify which entity in the sentence corresponds to an entity in the database, thereby creating a mapping. Using the entity names from the database, the LLM is prompted to reformulate the question. If a database query results in multiple rows for a selected column and item, a new question is generated using a greedy approach, prompting the user for more specific details until a conclusive answer is reached. Experiments show that this method exhibits notable enhancements in comparison to contemporary state-of-the-art techniques in mitigating language model inaccuracies.
StructGPT (Jiang et al., 2023b) is a general prompt framework to support LLMs reasoning on structured data (e.g., KG, Table, and Database). In general, the core of this framework is that they construct the specialized interfaces to collect relevant evidence from structured data (i.e., reading), and let LLMs concentrate on the reasoning task based on the collected information (i.e., reasoning). Specially, they propose an invoking-linearization-generation procedure to support LLMs in reasoning on the structured data with the help of the interfaces. By iterating this procedure with provided interfaces, our approach can gradually approach the target answers to a given query. Experiments conducted on three types of structured data, including KGQA, TableQA, and Text-to-SQL, show that StructGPT greatly improves the performance of LLMs, under the few-shot or zero-shot settings.
Baek et al. (2023) propose to inject the factual knowledge from knowledge graphs into (large) language models (up to GPT-3.5), by retrieving the relevant facts from knowledge graphs based on their textual similarities with the input question and then injecting them as the prompt of language models. This approach improves the performance of language models on knowledge graph question answering tasks (Sen et al., 2022) by up to 48% on average, compared to baselines without knowledge graphs.
5.3. Domain Factuality Enhanced LLMs
@@
Table 9. LLMs Enhanced for Domain-Specific Factuality. In the ‘domain’ column, we utilize the following abbreviations: healthcare/medicine (H), finance (F), law/legal (L), geoscience/environment (G), education (E), food testing (FT), and home renovation (HR).
Reference Do- main Base-Model Eval Task Eval Dataset Continual Pretrained? Continual SFT? Train From Scratch? External Knowledge HuatuoGPT (Zhang et al., 2023a) H Baichuan-7B, Ziya-LLaMA-13B QA cMedQA2, WebMedQA, Huatuo-26M Zhongjing (Yang et al., 2023c) H Ziya-LLaMA-13B QA CMtMedQA, huatuo-26M Wang et al. (2023f) H GPT-3.5-Turbo, LLaMA-2-13B QA MedQAUSMLE, MedQAMCMLE, MedMCQA MOLFORMER (Ross et al., 2022) H MOLFORMER Molecule properties prediction DISC-MedLLM (Bao et al., 2023) H Baichuan-13B QA CMB-Clin, CMD, CMID CohortGPT (Guan et al., 2023) H ChatGPT IU-RR, MIMIC-CXR DeID-GPT (Liu et al., 2023e) H GPT-4 Medical Text De-Identification ChatDoctor (Li et al., 2023b) H LLaMA QA BioMedLM (Venigalla et al., 2022) H GPT (2.7b) QA DoctorGLM (Xiong et al., 2023b) H ChatGLM-6B QA MedChatZH (Tan et al., 2023a) H Baichuan-7B QA C-Eval, MMLU BioGPT (Luo et al., 2022) H GPT-2 QA, DC, RE GeneGPT (Jin et al., 2023) H Codex QA GeneTuring Almanac (Zakka et al., 2023) H text-davinci-003 QA ClinicalQA MolXPT (Liu et al., 2023f) H GPT-2medium Molecular Property Prediction, Molecule-text translation LawGPT (Nguyen, 2023) L GPT3 Savelka et al. (2023) L GPT-4 Lawyer LLaMA (Huang et al., 2023c) L LLaMA CN Legal Tasks ChatLaw (Cui et al., 2023a) L Ziya-LLaMA-13B QA national judicial examination question EcomGPT (Li et al., 2023c) F BLOOMZ 4 major tasks 12 subtasks EcomInstruct BloombergGPT (Wu et al., 2023) F BLOOM Financial NLP (SA, BC, NER, NER+NED, QA) Financial Datasets K2 (Deng et al., 2023) G LLaMA-7B GeoBench HouYi (Bai et al., 2023) G ChatGLM-6B OceanGPT (Bi et al., 2023) G LLaMA-2-7B OceanBench GrammarGPT (Fan et al., 2023) E phoenix-inst-chat-7b Chinese Grammatical Error Correction ChatGPT- generated, Human- annotated FoodGPT (Qi et al., 2023) FT Chinese-LLaMA2-13B QA ChatHome (Wen et al., 2023) HR Baichuan-13B C-Eval, CMMLU, EvalHome
Table 9 lists the domain-factuality enhanced LLMs. Here, we include several domains, including healthcare/medicine (H), finance (F), law/legal (L), geoscience/environment (G), education (E), food testing (FT), and home renovation (HR).
@@
(2) Continual SFT: Another strategy for enhancing the factuality of AI models. In this approach, the model is fine-tuned using labeled or annotated data specific to the domain of interest. This fine-tuning process allows the model to learn and adapt to the nuances and specifics of the domain, improving its ability to provide accurate and contextually relevant information. It can be particularly useful in applications where access to domain-specific labeled data is available over time, such as in the case of legal databases, medical records, or financial reports (Bao et al., 2023; Li et al., 2023b).
@@
For each Domain-specific LLM, we list its respective enhancement methods, which are presented in Table 9.
5.4. Healthcare domain-enhanced LLMs
These LLMs have emerged as powerful tools in the medical field, offering a diverse range of capabilities. These models, such as CohortGPT(Guan et al., 2023), ChatDoctor(Li et al., 2023b), DeID-GPTLiu et al., 2023e, BioMedLMVenigalla et al., 2022, DoctorGLMXiong et al., 2023a, MedChatZHTan et al., 2023a, BioGPT(Luo et al., 2022), GeneGPT(Jin et al., 2023), Almanac(Zakka et al., 2023), and MolXPT(Liu et al., 2023f), harness the potential of LLMs to revolutionize healthcare. They are equipped with features like classifying unstructured medical text into disease labels, improving performance with knowledge graphs and sample selection strategies, fine-tuning on large datasets of patient-doctor dialogues, enabling automatic medical text de-identification, excelling in medical question-answering tasks, handling traditional Chinese medical question-answering, and outperforming in various biomedical NLP tasks. Some models interact with web APIs for genomics questions, while others specialize in clinical guidelines and treatment recommendations. These LLMs not only demonstrate state-of-the-art performance on healthcare domain but also emphasize the importance of domain-specific training and evaluation, showcasing their potential in transforming healthcare and clinical decision-making.